# Global thermodynamic manifold for conservative control of stochastic systems

### v2

@@ Acknowledgements.
Appendix A Full-control condition
A set of control parameters $\bm{u}$ is complete when it can manipulate the energy landscape $\bm{V}(\bm{u})$ up to a global additive constant. For $|\Omega|=n$, this requires $n-1$ parameters. Formally, the completeness condition may be stated as follows: for all $\bm{V}^{\prime}\in\mathbb{R}^{m}$, there exists some $\bm{u}\in\mathbb{R}^{n-1}$ and $c\in\mathbb{R}$ such that
$\bm{V}^{\prime}=\bm{V}(\bm{u})+c\bm{1}\ ,\ \ \forall\,\bm{V}^{\prime}\in\mathbb{R}^{n}\ .$ (78)
Assume that the first derivatives of $\bm{V}(\bm{u})$ in the parameters $u^{i}$ are continuous. We show that (78) is equivalent to the condition that $\left\langle\bm{f}\right\rangle_{\bm{p}(t)}=-\mathbb{J}\bm{p}(t)$ has inverse
@@ Appendix A Full-control condition
First we show that (78$\implies$(79). The condition (78) implies that $\mathbb{J}$ has full row rank, so the right inverse of the Jacobian for a complete control set exists, and satisfies
@@
To prove the converse (i.e., (79$\implies$(78)), the existence of a right inverse $\mathbb{J}^{\mathsf{R}}$ immediately implies that $\mathcal{N}(\mathbb{J})$ is one-dimensional. Left-multiplying (81) by $\mathbb{J}$ and applying (80a) gives
@@ Appendix D Spectral decomposition of the Drazin inverse
(((( $\mathcal{G}\circ\mathcal{G}^{\mathcal{D}})(x,y)$
@@
To see that properties (35a,b) hold with index $k=1$, note that $01\mathcal{G}^{\mathcal{D}}=\mathcal{G}^{\mathcal{D}}\psi_{0}=$ 0 by construction. Then
@@
Since the Drazin inverse is the unique operator satisfying properties (35a-c), $\mathcal{G}^{\mathcal{D}}$ is the Drazin inverse of $\mathcal{G}$.
Appendix E Proof of Equation (45)
@@ Appendix F Right inverse of the Ising full-control Jacobian
For $k\neq\ell$, we need only consider spin indices in either $\mathcal{C}(k)$ or $\mathcal{C}(\ell)$, since any $i\in\mathcal{C}(k)\cap\mathcal{C}(\ell)$ contributes a factor of one to the product. Denote this set by $\mathcal{C}(k)\oplus\mathcal{C}(\ell)$. For each state $x_{\alpha}$, there is a state $x_{\overline{\alpha}}$ obtained by inverting all spins. If $\mathcal{C}(k)\oplus\mathcal{C}(\ell)$ has an odd number of spins, then the $\alpha$th and $\overline{\alpha}$th terms in the sum cancel, so these entries will vanish. If $\mathcal{C}(k)\oplus\mathcal{C}(\ell)$ has an even number of spins, a similar cancellation occurs by inverting only the first spin in each (ordered) product. Therefore, $\left[\mathbb{J}\mathbb{J}^{\mathsf{T}}\right]_{k\ell}=2^{N}\delta_{k\ell}$, so
@@ Appendix G Control sets and numerical methods for the four-spin system
We computed minimum-work protocols for a collection of 200 control-parameter sets with between $1$ and $15$ parameters for the four-spin system. The endpoints for all protocols are $h(0)=-1$ and $h(\tau)=+1$ for global field $h\ (h_{0}=h_{1}=h_{2}=h_{3})$, $J_{\text{NN}}(0)=J_{\text{NN}}(\tau)=1$ for “nearest-neighbour” coupling $J_{\text{NN}}\ (J_{01}=J_{02}=J_{13}=J_{23})$, and all other interactions set to zero. Any control set capable of driving the system between these endpoints must have (at minimum) control over all of the fields $h_{i}$.
@@
×[$\ \ \ \ \ \times\left[\neg J,J_{\text{NN}},J_{\text{NNN}},(J_{\text{NN}},J_{\text{NNN}}),J_{ij}\right]$
×[$\ \ \ \ \ \times\left[\neg K,K,K_{=},K_{\times},K_{ijk}\right]$
×[$\ \ \ \ \ \times\left[\neg L,L\right]\ .$ (103)
The elements in the sets above are control subsets defined by constraints on the $1$-, $2$-, $3$-, and $4$-spin interactions. For example, referring to Table 1, the subset $h_{=}$ contains two parameters, $J_{\text{NNN}}$ and $L$ each contain one parameter, and $\neg K$ is an empty set, so the control-parameter set $(h_{=},J_{\text{NNN}},\neg K,L$) contains the four parameters

### v3

@@ VIII Summary and Outlook
A natural extension concerns higher-order corrections to the excess work. From the relation (27) and the operator decomposition (43), one may iteratively construct an asymptotic series of the excess work in powers of $\tau^{-1}$, as shown in [46] for discrete systems. Future research on this subject could address the convergence properties of the formal series and optimal truncation.
@@ Acknowledgements.
Data availability
The data that support the findings of this article are openly available [68].
Appendix A Completeness condition
A set of control parameters $\bm{u}$ is complete when it can manipulate the energy landscape $\bm{V}(\bm{u})$ up to a global additive constant. For $|\Omega|=n$, this requires n1n-1 parameters. We show that when the hypersurface $\bm{V}(\bm{u})$ is perpendicular to the null direction $\bm{1}$, the completeness condition Eq. (31) is equivalent to the statement that $\left\langle\bm{f}\right\rangle_{\bm{p}(t)}=-\mathbb{J}\bm{p}(t)$ has inverse
$\bm{p}(t)=-\mathbb{J}^{\mathsf{R}}\left\langle\bm{f}(t)\right\rangle+\frac{1}{n}\bm{1}\$(79)
@@ Appendix A Completeness condition
[…]
@@
[…]
@@ Appendix D Spectral decomposition of the Drazin inverse
$($ $\mathcal{G}\circ\mathcal{G}^{\mathcal{D}})(x,y)$
@@
To see that properties (36a,b) hold with index $k=1$, note that $1\mathcal{G}^{\mathcal{D}}=\mathcal{G}^{\mathcal{D}}\psi_{0}=0$ by construction. Then
@@
Since the Drazin inverse is the unique operator satisfying properties (36a)–(36c), $\mathcal{G}^{\mathcal{D}}$ is the Drazin inverse of $\mathcal{G}$.
Appendix E Proof of Eq. (46)
@@ Appendix F Right inverse of the Ising full-control Jacobian
[…]
@@ Appendix G Control sets and numerical methods for the four-spin system
We computed minimum-work protocols for a collection of 200 control-parameter sets with between $1$ and $15$ parameters for the four-spin system. The endpoints for all protocols are h(0)=1h(0)=-1 and $h(\tau)=+1$ for global field $h\ (h_{0}=h_{1}=h_{2}=h_{3})$, $J_{\text{NN}}(0)=J_{\text{NN}}(\tau)=1$ for “nearest-neighbor” coupling $J_{\text{NN}}\ (J_{01}=J_{02}=J_{13}=J_{23})$, and all other interactions set to zero. Any control set capable of driving the system between these endpoints must have (at minimum) control over all of the fields $h_{i}$.
@@
$\ \ \ \ \ \times\left[\neg J,J_{\text{NN}},J_{\text{NNN}},(J_{\text{NN}},J_{\text{NNN}}),J_{ij}\right]$
$\ \ \ \ \ \times\left[\neg K,K,K_{=},K_{\times},K_{ijk}\right]$
$\ \ \ \ \ \times\left[\neg L,L\right]\ .$ (103)
The elements in the sets above are control subsets defined by constraints on the $1$-, 22-, 33-, and $4$-spin interactions. For example, referring to Table 1, the subset $h_{=}$ contains two parameters, $J_{\text{NNN}}$ and $L$ each contain one parameter, and $\neg K$ is an empty set, so the control-parameter set $(h_{=},J_{\text{NNN}},\neg K,L$) contains the four parameters